A body is dropped from a certain height and takes 6 second to reach the ground.How much time does it take to cover one-ninth the distance starting from top?

Grade:9

1 Answers

Susmita
5 years ago
h=ut+ \frac{1}{2}gt^2
initial speed u=0.
t=6s
acceleration g=9.81 m/s2=10m/s2 approx.
Therefore , height h= \frac{1}{2}*10*6^2=5*36 m
Now one nineth the distance  is 
H=\frac{h}{9}= \frac{1}{9}*5*36=5*4 m=20m
Using the same formula
t= \sqrt[]{2h/g}=\sqrt[]{\frac{2*20}{10}}=2s